OpenStudy (anonymous):

Simplify −0.5(6.5). −3.5 −3.25 3.25 3.5

pooja195 (pooja195):

ignore the sign for right now -0.5 x 6.5=?

OpenStudy (anonymous):

im bad a multiplying .....

OpenStudy (photon336):

\[\frac{ -1 }{ 2 }*\frac{ 6.5 }{ 1}= (\frac{ -6.5 }{ 2}) = -3.25 \]

OpenStudy (anonymous):

r u allowed to give the answer like that o, o

OpenStudy (anonymous):

either way thx

OpenStudy (photon336):

you could probably just rewrite it like 6.5/1 and then when you multiply fractions, you multiply both the numerator and denominator

OpenStudy (anonymous):

nailed the test thx guys

OpenStudy (anonymous):

ill give medal to pooja cause she actually tryed to explain.....

pooja195 (pooja195):